OpenStudy (anonymous):

how would i solve 4x^2-1=8x by completing the square and have it in the form x=a+-radical b over c, where a, b and c have no common factors??

OpenStudy (anonymous):

idk that one srrrry

OpenStudy (anonymous):

ok so first convert it into a quadratic in standard form so you get: \[4x^2-8x-1=0\] (i subtracted 8x from both sides) now divide the whole equation by 4 so you get:\[x^2-2x-\frac{ 1 }{ 4 }=0\] Now add 1/4 to both sides: \[x^2-2x=\frac{ 1 }{ 4 }\] now add half of the b-term squared to both sides (-1)^2 so 1:\[x^2-2x+1=\frac{ 5 }{ 4 }\] (i added 1 to both sides). Now you can rewrite that as (x+half of the b-term)^2:\[(x-1)^2=\frac{ 5 }{ 4 }\] now take the plus/minus square root so:\[\pm \sqrt{(x-1)^2}=\pm\sqrt{\frac{ 5 }{ 4 }}\] so if you simplify that:\[x-1=\pm \frac{ \sqrt{5} }{ 2 }\] now add 1 to both sides so:\[x=\pm \frac{ \sqrt{5} }{ 2 }+1\] as you can see 2, sqrt(5), and 1 have no common factors but 1 so thats the answer! THE ANSWER IS: \[x=\pm \frac{ \sqrt{5} }{ 2 }+1\]

OpenStudy (anonymous):

\[\pm \] means positive or negative and read my work carefully.